Opened 2 weeks ago

Last modified 21 hours ago

#627 new flight processing

Fenix Calibration, 2018

Reported by: asm Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords: calibration, fenix
Cc: Other processors:

Description

This ticket records the calibration of the Fenix sensor. The Fenix sensor was repaired by Specim and the detector array was replaced with a new one so timeline comparisons should not be made as it is essentially a new sensor. Specim also performed their own calibration in February 2018

Data location:
~arsf/arsf_data/2018/misc/calibration_april/original_data

Change History (2)

comment:1 Changed 2 weeks ago by asm

-The 2018 Calibration took 4 days: 16-19 April 2018.

-The fenix sensor needed a replacement of the array so the new wavelengths are expected to be different

-Specim performed its own calibration on Feb 2018. On a first look the new hdr files by Specim looks to be in accordance with our wavelength calibration

-Details recorded are within the spreadsheet: ./data_details/recorded_at_BAS_lab/Capture_Spreadsheet_2018-04.ods

-In short, the calibration went as follows:

Day1: Fenix wavelength cal+radiometric cal (no dark frames for 1 spectral lamp and some radiometric files). Also some Owl validation data with the black body.
Day2: Monochromator tests for different ranges of wavelengths in the SWIR +1 in the VNIR (no Fenix calibration data)
Day3: Fenix wavelength cal+radiometric cal.
Day4: Fenix wavelength cal+radiometric cal. Also some Owl validation data with the black body.

comment:2 Changed 21 hours ago by asm

Wavelength accuracy.

After playing quite a lot with the data I think I am happy now. I have decided not to use the Mylar filter that provides some features in the SWIR region because its accuracy is 1nm and we need to subtract the FWHMs of the same to give an approximation. We did have good data but the most precise is from day 1. Just a quick summary:

-Day 1: Best fit 33 wavelengths used, 6 of them for the SWIR. The offset is 0.076660 for the VNIR and -1.199219
The Oxygen lamp could not be used for d1 because it did not have dark frames but even though is not providing any usable feature for the other days as it is too weak.
-Day 3: We did not use the black cloth so the data quality is good but the other is even better. 35 wavelengths but 2 of them failed. We had some an extra wavelengths for 404.7 nm (Hg) but this region has very low sensitivity so should not be used. The offset for this day was: 0.370117 and -0.218750
-Day 4: Put back the black cloth, also good data with 30 wavelengths used but only 3 on the SWIR. Somehow we could not use the wavelenghts of the Hg-Ar that provides good anchors on the SWIR. Again the scripts could use the 404.7 nm peak which I think it should not be used. The offset for this case is: 0.190430 and 0.200000

As you can see, offsets are small in this season and all on the same region so the calibration made by Specim was nice. For all the above, we will use the wavelength calibration obtained from the data in the day 1. Right now is on my scratch only but I will move it so a safe place once done the full calibration procedure.

Note: See TracTickets for help on using tickets.